OpenStudy (anonymous):

Please help! Solve for x if 4+ix=2x+(y-x)i. x= a. 0 b. 1 c. 2 d. 4

OpenStudy (anonymous):

Compex numbers, right? Well every complex number is actually a combination of two different things, a real number and an imaginary number. So when two complex numbers equal each other, both their real and imaginary numbers must equal each other.\[\large \text{If}~~a+ib=c+id,~~\text{then}~~a=c~~~\text{and}~~~b=d\]So\[\large 4=2x~~\text{and}~~x=y-x\]But you only need one of these equations to solve it! Do you understand now?

OpenStudy (anonymous):

Oh o.k. Thanks! Is the answer 2? Just for confirmation?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

How would I find y?

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

Plug x into the second equation and solve. Answer in white letters below. \[\color{white}{y=4}\]

OpenStudy (anonymous):

Thank you!